# Sample complexity of optimal transport barycenters with discrete support

### v2

@@ Appendix B Boundedness of the dual variables
Let $\mu\in\mathcal{M}_{1}(\mathbb{B}_{R})$ for some $R>0$ and $\mathbb{D}=W_{p}^{p},\>\text{or}\>W_{\epsilon,p}^{p}$. Then for any $(Y,\pi)\in(\mathbb{B}_{R}^{N}\setminus D_{N})\times\mathrm{int}(\Delta_{N})$, we have the following
$W_{p}^{p}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)=\underset{\mid w_{i}\mid\leq 2pR^{p}}{\underset{w\in\mathbb{R}^{N}}{\max}}\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,N}\{\|x-y_{i}\|^{p}-w_{i}\}d\mu(x)+\sum_{i=1}^{N}\pi_{i}w_{i},$
@@
$W_{\epsilon,p}^{p}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)=\underset{\mid w_{i}\mid\leq 4p(2R)^{p}}{\underset{w\in\mathbb{R}^{N}}{\max}}-\epsilon\int_{\mathbb{R}^{d}}\log\left(\sum_{i=1}^{N}\pi_{i}\exp\left(\frac{w_{i}-\|x-y_{i}\|^{p}}{\epsilon}\right)\right)d\mu(x)+\sum_{i=1}^{N}\pi_{i}w_{i}-\epsilon.$
@@
As explained in Remark (B.2), we may select a choice of dual variables $w\in\mathbb{R}^{N}$ optimal in (B.1) and verifying $w_{N}=0$. From (book_santambrogio, Proposition 5.11) and (book_santambrogio, Remark 1.13) that follows, we may also select a c-convave function $w^{c}$ which verifies for all $i\in[\![1,N]\!]$: $w_{i}=(w^{c}(y_{i}))^{c}$. We then have
@@
$\leq pR^{p-1}\|z_{i}-y_{i}\|$ (B.3)
$\leq 2pR^{p},$
where inequality (B.2) comes from Lemma D.2 and (B.3) from Lemma D.1. Using that $w_{N}=0$, we get the result, namely, the dual variables for the $p$-Wasserstein distances are bounded, and $|w_{i}|\leq 2pR^{p}$.
@@
Let $(Y,\pi)\in(\mathbb{B}_{R}^{N}\setminus D_{N})\times\mathrm{int}(\Delta_{N})$. Let also $u\in\mathcal{C}_{b}(\mathbb{B}_{R})$ be an optimal solution in the dual formulation of the entropic $p$-Wasserstein distance between $\mu$ and $\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}$ given in (A.3). By (genevay2019sample, Proposition 1), $u$ is $\kappa$-Lipshitz, where $\kappa$ is the Lipschitz constant of $(a,b)\in\mathbb{B}_{R}\times\mathbb{B}_{R}\mapsto\|a-b\|^{p}$. Then since $\operatorname{Supp}(\mu)\subset\mathbb{B}_{R}$, we have by Lemma D.1 that $\kappa$ is upper bounded by $p\sqrt{2}(2R)^{p-1}$. Let $w$ be a solution of (A.4) for which we can suppose $w_{N}=0$ (Remark (B.2)). As in genevay2016stochastic, one can show with a first order condition on (A.3) that for all $y$ a coordinate in $Y$:
$w(y)=-\epsilon\log\left(\int_{\mathbb{R}^{d}}\exp\left(\frac{u(x)-\|x-y\|^{p}}{\epsilon}\right)d\mu(x)\right).$
We thus have for any $y_{1},y_{2}$ two distinct points in the point cloud $Y$,
$|w(y_{1})-w(y_{2})|$ $=\left|\epsilon\log\left(\frac{\int_{\mathbb{R}^{d}}\exp\left(\frac{u(x)-\|x-y_{1}\|^{p}}{\epsilon}\right)d\mu(x)}{\int_{\mathbb{R}^{d}}\exp\left(\frac{u(x)-\|x-y_{2}\|^{p}}{\epsilon}\right)d\mu(x)}\right)\right|$
$\leq\left|\epsilon\log\left(\frac{\int_{\mathbb{R}^{d}}\exp\left(\frac{u(\overline{x}_{1})-\|\overline{x}_{1}-y_{1}\|^{p}}{\epsilon}\right)d\mu(x)}{\int_{\mathbb{R}^{d}}\exp\left(\frac{u(\underline{x}_{2})-\|\underline{x}_{2}-y_{2}\|^{p}}{\epsilon}\right)d\mu(x)}\right)\right|$
$=|u(\overline{x}_{1})-\|\overline{x}_{1}-y_{2}\|^{p}-(u(\underline{x}_{2})-\|\underline{x}_{2}-y_{1}\|^{p})|$
$\leq|u(\overline{x}_{1})-u(\underline{x}_{2})|+|\ \|\overline{x}_{1}-y_{2}\|^{p}-\|\underline{x}_{2}-y_{1}\|^{p}|$
$\leq p\sqrt{2}(2R)^{p-1}\|\overline{x}_{1}-\underline{x}_{2}\|+\|\overline{x}_{1}-y_{2}\|^{p}+\|\underline{x}_{2}-y_{1}\|^{p}$ (B.4)
$\leq p\sqrt{2}(2R)^{p}+2(2R)^{p}$
$\leq 4p(2R)^{p},$
@@
where $\overline{x}_{1}$ and $\underline{x}_{2}$ denote respectively a maximizer of $x\mapsto\exp\left(\frac{u(x)-\|x-y_{1}\|^{p}}{\epsilon}\right)$ and a minimizer of $x\mapsto\exp\left(\frac{u(x)-\|x-y_{2}\|^{p}}{\epsilon}\right)$ over $\mathbb{B}_{R}$. Using that $w(y_{N})=0$, we get that for all $i\in[\![1,N]\!]$,
$|w(y_{i})|\leq 4p(2R)^{p}.$
Appendix C Proofs of the main theorems
In this section, we prove the main result in Theorem 4.1, which is the sample complexity of optimal transport barycenters with sparse support. Throughout the section, $\mathbb{P}_{n}^{\ell}$ for $\ell\in[\![1,L]\!]$ denotes the empirical process
$\mathbb{P}_{n}^{\ell}:f\mapsto\mathbb{P}_{n}^{\ell}f:=\frac{1}{n}\sum_{i=1}^{n}f(X_{i}^{\ell})$
@@ Appendix C Proofs of the main theorems
for $n$ i.i.d random variables $X_{1}^{\ell},\ldots,X_{n}^{\ell}$ of law $\mu^{\ell}$. For any subset $\mathcal{F}$ of a normed space $(\Theta,\|\cdot\|)$ and $\tau>0$, we denote $\mathcal{N}(\tau,\mathcal{F},\|\cdot\|)$ the $\tau$-covering number of $\mathcal{F}$ with respect to $\|\cdot\|$. This quantity represents the minimum number of balls of radius $\tau$, with respect to the norm $\|\cdot\|$, needed to fully cover $\mathcal{F}$. We refer the reader to (sen2018gentle, Definition 2.2) for more details on the covering number. Finally, we also denote $\|\cdot\|_{\mathbb{L}^{2}(\mu)}$ the $\mathbb{L}^{2}$ norm with respect to the measure $\mu$ and $\overline{\text{B}_{d}(0,R)}$ the $d$-dimensional closed ball of radius $R$.
The proofs for Wasserstein barycenters (Section C.1), Sliced Wasserstein barycenter (Section C.2) and entropy regularized barycenters (Section C.3) all follow the same approach : we first derive the covering number for a class of functions associated to the dual formulation of the divergences $\mathbb{D}$, and then apply tools from empirical process theory to obtain a generalization error bound for the barycenters.
@@ C.1 Wasserstein barycenter
then for all $\tau>0$ and $\mu\in\mathcal{M}_{1}(\mathbb{B}_{R})$:
$\mathcal{N}(\tau,\mathcal{F}_{p},\|\cdot\|_{\mathbb{L}^{2}(\mu)})\leq\left(1+\frac{4pR^{p}}{\tau}\right)^{dN}\left(1+\frac{8pR^{p}}{\tau}\right)^{N}$
Moreover $\mathcal{F}_{p,\mathbb{Q}}$, the subset of functions of $\mathcal{F}_{p}$ whose parameters $(Y,w)$ are rationals, is dense in $\mathcal{F}_{p}$.
@@
where inequality (C.2) is due to Lemma D.2 and inequality (C.3) is given by Lemma D.1. Let us denote $T=pR^{p-1}$, and let $\tau>0$. We then consider $z^{1},\ldots,z^{n(\tau)}$ a $\frac{\tau}{2T}$-covering of $\mathbb{B}_{R}$ and $v^{1},\ldots v^{m(\tau)}$ a $\frac{\tau}{2}$-covering of $\overline{\text{B}_{1}(0,2pR^{p})}$. For any vector of indices $i=(i_{1},\ldots,i_{N})\in[\![1,n(\tau)]\!]^{N}$ and $k=(k_{1},\ldots,k_{N})\in[\![1,m(\tau)]\!]^{N}$, we define for all $x\in\mathbb{B}_{R}$ the function $f_{i,k}(x)=\underset{j=1,...,N}{\min}\|x-z^{i_{j}}\|^{p}-v^{k_{j}}$. Let $f:=f_{Y}^{w}\in\mathcal{F}_{p}$, then by (C.3) there exists $i,k$ such that
$\begin{split}|f(x)-f_{i,k}(x)|&\leq pR^{p-1}\max_{j=1,...,N}\|y_{j}-z^{i_{j}}\|+\max_{j=1,...,N}|w_{j}-v^{k_{j}}|\\ &\leq T\frac{\tau}{2T}+\frac{\tau}{2}=\tau,\end{split}$
so that $\|f-f_{i,k}\|_{\mathbb{L}^{2}(\mu)}\leq\tau$.
Since there exists $n(\tau)^{N}\times m(\tau)^{N}$ possibilities for such functions $f_{i,k}$, it follows from the covering construction that :
[…]
where we second inequality is due to (massart2007concentration, Lemma 4.14).
@@
Theorem C.2.
Let $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{B}_{R})$ for some $R>0$. For some integer $n$, let $\mu_{n}^{1},\ldots,\mu_{n}^{L}$ be empirical measures supported over $n$ i.i.d random variables of respective law $\mu^{\ell}$ for all $\ell\in[\![1,L]\!]$, then for $\mathbb{D}=W_{p}^{p}$ and some integer $N\geq 1$, we have
$\begin{split}&\mathbb{E}\left[\sup_{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}\left|\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{\ell},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\right]\\ &\leq C_{p,R}\sqrt{\frac{N(d+1)}{n}},\end{split}$
where $C_{p,R}=8\sqrt{2}\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2+\frac{32pR^{p}}{\tau}\right)}d\tau$.
Proof.
@@
$\begin{split}&\left|\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\\ &=\left|\frac{1}{L}\sum_{\ell=1}^{L}\left[\underset{\mid\tilde{w}_{i}\mid\leq K}{\underset{\tilde{w}\in\mathbb{R}^{M}}{\max}}\frac{1}{n}\sum_{k=1}^{n}\underset{i=1,\ldots,M}{\min}\{\|X_{k}^{\ell}-\tilde{y}_{i}\|^{p}-\tilde{w}_{i}\}+\sum_{i=1}^{M}\tilde{\pi}_{i}\tilde{w}_{i}\right.\right.\\ &\left.\left.-\underset{\mid\tilde{w}_{i}\mid\leq K}{\underset{\tilde{w}\in\mathbb{R}^{M}}{\max}}\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,M}{\min}\{\|x-\tilde{y}_{i}\|^{p}-\tilde{w}_{i}\}d\mu^{\ell}(x)+\sum_{i=1}^{M}\tilde{\pi}_{i}\tilde{w}_{i}\right]\right|\\ &\leq\frac{1}{L}\sum_{\ell=1}^{L}\underset{\mid\tilde{w}_{i}\mid\leq K}{\underset{\tilde{w}\in\mathbb{R}^{M}}{\max}}\left|\frac{1}{n}\sum_{k=1}^{n}\underset{i=1,\ldots,M}{\min}\{\|X_{k}^{\ell}-\tilde{y_{i}}\|^{p}-\tilde{w}_{i}\}-\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,M}{\min}\{\|x-\tilde{y}_{i}\|^{p}-\tilde{w}_{i}\}d\mu^{\ell}(x)\right|\\ &=\frac{1}{L}\sum_{\ell=1}^{L}\underset{w_{i}=w_{j}\>\text{if}\>y_{i}=y_{j}}{\underset{\mid w_{i}\mid\leq K}{\underset{w\in\mathbb{R}^{N}}{\max}}}\left|\frac{1}{n}\sum_{k=1}^{n}\underset{i=1,\ldots,N}{\min}\{\|X_{k}^{\ell}-y_{i}\|^{p}-w_{i}\}-\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,N}{\min}\{\|x-y_{i}\|^{p}-w_{i}\}d\mu^{\ell}(x)\right|.\end{split}$
where the inequality is verified thanks to by Lemma D.2. The last equality is due to the fact that for all $Y\in\mathbb{B}_{R}^{N}$ and $x\in\mathbb{R}^{d}$, we have
$\min_{k=1,\ldots,M}\|x-y_{k}\|^{p}-w_{k}=\min_{k=1,\ldots,N}\|x-y_{k}\|^{p}-w_{k}$
if $w_{i}=w_{j}$ whenever $y_{i}=y_{j}$. Notice that the dependency in the vector of weights $\pi$ has disappeared.
For any $(x,Y,w)\in\mathbb{B}_{R}\times\mathbb{B}_{R}^{N}\times[-K,K]^{N}$, let $f_{Y}^{w}(x)=\underset{i=1,\ldots,N}{\min}\|x-y_{i}\|^{p}-w_{i}$. We then have
$\begin{split}&\underset{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}{\sup}\left|\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\\ &\leq\frac{1}{L}\sum_{\ell=1}^{L}\underset{(Y,w)\in\mathbb{B}_{R}^{N}\times[-K,K]^{N}}{\sup}\left|\mathbb{P}_{n}^{\ell}\left(f_{Y}^{w}\right)-\mathbb{E}_{X\sim\mu^{\ell}}\left[f_{Y}^{w}(X)\right]\right|\\ &=\frac{1}{L}\sum_{\ell=1}^{L}\underset{f\in\mathcal{F}_{p}}{\sup}|\mathbb{P}_{n}^{\ell}(f)-\mathbb{E}_{X\sim\mu^{\ell}}[f(X)]|\end{split}$
@@
which gives $\sqrt{\sup_{f\in\mathcal{F}}\mathbb{P}_{n}f^{2}}\leq 2(2R)^{p}$. By Lemma C.1, $\mathcal{F}_{p}$ is dense in $\mathcal{F}_{p,\mathbb{Q}}$, and we can then apply (gine2021mathematical, Theorem 3.5.1) on empirical processes and covering number. Therefore, denoting $\omega:=\underset{f\in\mathcal{F}}{\sup}\mathbb{P}_{n}f^{2}$, we get for any $\ell\in[\![1,L]\!]$ :
$\begin{split}\mathbb{E}\left[\sqrt{n}\underset{f\in\mathcal{F}_{p}}{\sup}|\mathbb{P}_{n}^{\ell}(f)-\mathbb{E}_{X\sim\mu^{\ell}}[f(X)]|\right]&\leq 8\sqrt{2}\mathbb{E}\left[\int_{0}^{\omega}\sqrt{\log\left(2\mathcal{N}(\tau/2,\mathcal{F}_{p},\|\cdot\|_{\mathbb{L}^{2}(\mu_{n}^{\ell})})\right)}d\tau\right]\\ &\leq 8\sqrt{2}\mathbb{E}\left[\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2\mathcal{N}(\tau/2,\mathcal{F}_{p},\|\cdot\|_{\mathbb{L}^{2}(\mu_{n}^{\ell})})\right)}d\tau\right].\end{split}$
From Lemma C.1 applied to the measure $\mu_{n}^{\ell}$, we get
$\begin{split}\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2\mathcal{N}(\tau/2,\mathcal{F}_{p},\|\cdot\|_{\mathbb{L}^{2}(\mu_{n}^{\ell})})\right)}d\tau&\leq\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2\left(1+\frac{8pR^{p}}{\tau}\right)^{dN}\left(1+\frac{16pR^{p}}{\tau}\right)^{N}\right)}d\tau\\ &\leq\int_{0}^{3(2R)^{p}}\sqrt{\log\left(2\left(1+\frac{16pR^{p}}{\tau}\right)^{N(d+1)}\right)}d\tau\\ &\leq\int_{0}^{2(2R)^{p}}\sqrt{\log\left(\left(2+\frac{32pR^{p}}{\tau}\right)^{N(d+1)}\right)}d\tau\\ &=\sqrt{N(d+1)}\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2+\frac{32pR^{p}}{\tau}\right)}d\tau.\end{split}$

### v3

@@ Appendix B Boundedness of the dual variables
Let $\mu\in\mathcal{M}_{1}(\mathbb{B}_{R})$ for some $R>0$ and $\mathbb{D}=W_{p}^{p},\>\text{or}\>W_{\epsilon,p}^{p}$. Then for any $(Y,\pi)\in(\mathbb{B}_{R}^{N}\setminus D_{N})\times\mathrm{int}(\Delta_{N})$, we have
$W_{p}^{p}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)=\underset{\mid w_{i}\mid\leq K_{1}}{\underset{w\in\mathbb{R}^{N}}{\max}}\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,N}\{\|x-y_{i}\|^{p}-w_{i}\}d\mu(x)+\sum_{i=1}^{N}\pi_{i}w_{i},$
@@
$W_{\epsilon,p}^{p}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)=\underset{\mid w_{i}\mid\leq K_{2}}{\underset{w\in\mathbb{R}^{N}}{\max}}-\epsilon\int_{\mathbb{R}^{d}}\log\left(\sum_{i=1}^{N}\pi_{i}\exp\left[\frac{w_{i}-\|x-y_{i}\|^{p}}{\epsilon}\right]\right)d\mu(x)+\sum_{i=1}^{N}\pi_{i}w_{i}-\epsilon$
where $K_{1}=2pR^{p}$ and $K_{2}=4p(2R)^{p}$.
@@
As discussed in Remark (B.2), we may select a choice of dual variables $w\in\mathbb{R}^{N}$ optimal in (B.1) and verifying $w_{N}=0$. From (book_santambrogio, Proposition 5.11) and (book_santambrogio, Remark 1.13) that follows, we may also select a c-convave function $w^{c}$ which verifies for all $i\in[\![1,N]\!]$: $w_{i}=(w^{c}(y_{i}))^{c}$. We then have
@@
$\leq pR^{p-1}\|z_{i}-y_{i}\|\leq 2pR^{p},$ (B.3)
where inequalities (B.2) and (B.3) come from Lemmas D.2 and D.1. Using that $w_{N}=0$, we get the result: the dual variables for the $p$-Wasserstein distances are bounded, and $|w_{i}|\leq 2pR^{p}$.
@@
Let $(Y,\pi)\in(\mathbb{B}_{R}^{N}\setminus D_{N})\times\mathrm{int}(\Delta_{N})$. Let also $u\in\mathcal{C}_{b}(\mathbb{B}_{R})$ be an optimal solution in the dual (A.3) formulation of the entropic $p$-Wasserstein distance between $\mu$ and $\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}$. By (genevay2019sample, Proposition 1), $u$ is $\kappa$-Lipshitz, where $\kappa$ is the Lipschitz constant of $(a,b)\in\mathbb{B}_{R}^{2}\mapsto\|a-b\|^{p}$. Then since $\operatorname{Supp}(\mu)\subset\mathbb{B}_{R}$, we have by Lemma D.1 that $\kappa$ is upper bounded by $p\sqrt{2}(2R)^{p-1}$. Let $w$ be a solution of (A.4) for which we can suppose $w_{N}=0$ (Remark (B.2)). As in genevay2016stochastic, one can show with a first order condition on (A.3) that for all $y$ a coordinate in $Y$: $w(y)=-\epsilon\log\left(\int_{\mathbb{R}^{d}}\exp\left(\frac{u(x)-\|x-y\|^{p}}{\epsilon}\right)d\mu(x)\right).$ We thus have for all $i\neq N$:
$|w(y_{i})-w(y_{N})|$ $=\left|\epsilon\log\left(\frac{\int_{\mathbb{R}^{d}}\exp\left(\frac{u(x)-\|x-y_{i}\|^{p}}{\epsilon}\right)d\mu(x)}{\int_{\mathbb{R}^{d}}\exp\left(\frac{u(x)-\|x-y_{N}\|^{p}}{\epsilon}\right)d\mu(x)}\right)\right|$
$\leq\left|\epsilon\log\left(\frac{\int_{\mathbb{R}^{d}}\exp\left(\frac{u(\overline{x}_{i})-\|\overline{x}_{i}-y_{i}\|^{p}}{\epsilon}\right)d\mu(x)}{\int_{\mathbb{R}^{d}}\exp\left(\frac{u(\underline{x}_{N})-\|\underline{x}_{N}-y_{N}\|^{p}}{\epsilon}\right)d\mu(x)}\right)\right|$
$=|u(\overline{x}_{i})-\|\overline{x}_{i}-y_{i}\|^{p}-(u(\underline{x}_{N})-\|\underline{x}_{N}-y_{N}\|^{p})|$
$\leq|u(\overline{x}_{i})-u(\underline{x}_{N})|+|\ \|\overline{x}_{i}-y_{i}\|^{p}-\|\underline{x}_{N}-y_{N}\|^{p}|$
$\leq p\sqrt{2}(2R)^{p-1}\|\overline{x}_{i}-\underline{x}_{N}\|+\|\overline{x}_{i}-y_{i}\|^{p}+\|\underline{x}_{N}-y_{N}\|^{p}$ (B.4)
$\leq p\sqrt{2}(2R)^{p}+2(2R)^{p}\leq 4p(2R)^{p},$
@@
where $\overline{x}_{i}$ and $\underline{x}_{N}$ denote respectively a maximizer of $x\mapsto\exp\left(\frac{u(x)-\|x-y_{i}\|^{p}}{\epsilon}\right)$ and a minimizer of $x\mapsto\exp\left(\frac{u(x)-\|x-y_{N}\|^{p}}{\epsilon}\right)$ over $\mathbb{B}_{R}$. Using that $w(y_{N})=0$ then yields the result.
Appendix C Proofs of the main results
@@
In this section, we prove the sample complexity of optimal transport barycenters with sparse support in Theorem 3.1. For $\ell\in[\![1,L]\!]$, we denote $\mathbb{P}_{n}^{\ell}$ the empirical process $f\mapsto\mathbb{P}_{n}^{\ell}f:=\frac{1}{n}\sum_{i=1}^{n}f(X_{i}^{\ell})$ for $n$ i.i.d random variables $X_{1}^{\ell},\ldots,X_{n}^{\ell}$ of law $\mu^{\ell}$. For any subset $\mathcal{F}$ of a normed space $(\Theta,\|\cdot\|)$ and $\tau>0$, we denote $\mathcal{N}(\tau,\mathcal{F},\|\cdot\|)$ the $\tau$-covering number of $\mathcal{F}$ with respect to $\|\cdot\|$, which is the minimum number of balls of radius $\tau$, with respect to the norm $\|\cdot\|$, needed to cover $\mathcal{F}$. We refer the reader to (sen2018gentle, Definition 2.2) for more details on the covering number. Finally, we also denote $\|\cdot\|_{\mathbb{L}^{2}(\mu)}$ the $\mathbb{L}^{2}$ norm with respect to the measure $\mu$ and $\overline{\text{B}_{d}(0,R)}$ the $d$-dimensional closed ball of radius $R$. The proofs for Wasserstein (Section C.1), Sliced Wasserstein (Section C.2) and Sinkhorn (Section C.3) barycenters all follow the same approach : we first derive the covering number for a class of functions associated to the dual formulation of the divergences $\mathbb{D}$, and then apply tools from empirical process theory to obtain a generalization error bound for the barycenters.
@@ C.1 Wasserstein barycenter
then for all $\tau>0$ and $\mu\in\mathcal{M}_{1}(\mathbb{B}_{R})$, $\mathcal{N}(\tau,\mathcal{F}_{p},\|\cdot\|_{\mathbb{L}^{2}(\mu)})\leq\left(1+\frac{4pR^{p}}{\tau}\right)^{dN}\left(1+\frac{8pR^{p}}{\tau}\right)^{N}$. Also, the subset $\mathcal{F}_{p,\mathbb{Q}}$ of functions of $\mathcal{F}_{p}$ with rationals parameters $(Y,w)$, is dense in $\mathcal{F}_{p}$.
@@
where inequalities (C.2) and (C.3) are due to Lemmas D.2 and D.1. Let us denote $T=pR^{p-1}$, and let $\tau>0$. We then consider $z^{1},\ldots,z^{n(\tau)}$ a $\frac{\tau}{2T}$-covering of $\mathbb{B}_{R}$ and $v^{1},\ldots v^{m(\tau)}$ a $\frac{\tau}{2}$-covering of $\overline{\text{B}_{1}(0,2pR^{p})}$. For any vector of indices $i=(i_{1},\ldots,i_{N})\in[\![1,n(\tau)]\!]^{N}$ and $k=(k_{1},\ldots,k_{N})\in[\![1,m(\tau)]\!]^{N}$, we define for all $x\in\mathbb{B}_{R}$ the function $f_{i,k}(x)=\underset{j=1,...,N}{\min}\|x-z^{i_{j}}\|^{p}-v^{k_{j}}$. Let $f:=f_{Y}^{w}\in\mathcal{F}_{p}$, then by (C.3) there exists $i,k$ such that
$|f(x)-f_{i,k}(x)|\leq pR^{p-1}\max_{j=1,...,N}\|y_{j}-z^{i_{j}}\|+\max_{j=1,...,N}|w_{j}-v^{k_{j}}|\leq T\frac{\tau}{2T}+\frac{\tau}{2}=\tau,$
so that $\|f-f_{i,k}\|_{\mathbb{L}^{2}(\mu)}\leq\tau$. Since there exists $n(\tau)^{N}\times m(\tau)^{N}$ possibilities for such functions $f_{i,k}$, it follows from the covering construction that :
$\begin{split}\mathcal{N}(\tau,\mathcal{F}_{p},\|\cdot\|_{\mathbb{L}^{2}(\mu)})&\leq\mathcal{N}\left(\frac{\tau}{2T},\overline{\text{B}_{d}(0,R)},\|\cdot\|\right)^{N}\mathcal{N}\left(\frac{\tau}{2},\overline{\text{B}_{1}(0,2pR^{p})},\mid\cdot\mid\right)^{N}\\ &\leq\left(1+\frac{4RT}{\tau}\right)^{dN}\left(1+\frac{8pR^{p}}{\tau}\right)^{N}=\left(1+\frac{4pR^{p}}{\tau}\right)^{dN}\left(1+\frac{8pR^{p}}{\tau}\right)^{N}\end{split}$
where the second inequality is due to (massart2007concentration, Lemma 4.14).
@@
Proof of Theorem 3.1 for $\mathbb{D}=W_{p}^{p}$.
@@
$\begin{split}&\text{Gen}_{n}(Y,\pi):=\left|\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\\ &\leq\frac{1}{L}\sum_{\ell=1}^{L}\underset{\mid\tilde{w}_{i}\mid\leq K}{\underset{\tilde{w}\in\mathbb{R}^{M}}{\max}}\left|\frac{1}{n}\sum_{k=1}^{n}\underset{i=1,\ldots,M}{\min}\{\|X_{k}^{\ell}-\tilde{y_{i}}\|^{p}-\tilde{w}_{i}\}-\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,M}{\min}\{\|x-\tilde{y}_{i}\|^{p}-\tilde{w}_{i}\}d\mu^{\ell}(x)\right|\\ &=\frac{1}{L}\sum_{\ell=1}^{L}\underset{w_{i}=w_{j}\>\text{if}\>y_{i}=y_{j}}{\underset{\mid w_{i}\mid\leq K}{\underset{w\in\mathbb{R}^{N}}{\max}}}\left|\frac{1}{n}\sum_{k=1}^{n}\underset{i=1,\ldots,N}{\min}\{\|X_{k}^{\ell}-y_{i}\|^{p}-w_{i}\}-\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,N}{\min}\{\|x-y_{i}\|^{p}-w_{i}\}d\mu^{\ell}(x)\right|.\end{split}$
where the inequality is due to Lemma D.2. The last equality is due to the fact that for all $Y\in\mathbb{B}_{R}^{N}$ and $x\in\mathbb{R}^{d}$, we have $\min_{k=1,\ldots,M}\|x-y_{k}\|^{p}-w_{k}=\min_{k=1,\ldots,N}\|x-y_{k}\|^{p}-w_{k}$ if $w_{i}=w_{j}$ whenever $y_{i}=y_{j}$. Notice that the dependency in the vector of weights $\pi$ has disappeared. For any $(x,Y,w)\in\mathbb{B}_{R}\times\mathbb{B}_{R}^{N}\times[-K,K]^{N}$, let $f_{Y}^{w}(x)=\underset{i=1,\ldots,N}{\min}\|x-y_{i}\|^{p}-w_{i}$. We then have
$\begin{split}\underset{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}{\sup}\text{Gen}_{n}(Y,\pi)&\leq\frac{1}{L}\sum_{\ell=1}^{L}\underset{(Y,w)\in\mathbb{B}_{R}^{N}\times[-K,K]^{N}}{\sup}\left|\mathbb{P}_{n}^{\ell}\left(f_{Y}^{w}\right)-\mathbb{E}_{X\sim\mu^{\ell}}\left[f_{Y}^{w}(X)\right]\right|\\ &=\frac{1}{L}\sum_{\ell=1}^{L}\underset{f\in\mathcal{F}_{p}}{\sup}|\mathbb{P}_{n}^{\ell}(f)-\mathbb{E}_{X\sim\mu^{\ell}}[f(X)]|\end{split}$
@@
which gives $\sqrt{\sup_{f\in\mathcal{F}}\mathbb{P}_{n}f^{2}}\leq 2(2R)^{p}$. By Lemma C.1, $\mathcal{F}_{p}$ is dense in $\mathcal{F}_{p,\mathbb{Q}}$, and we can then apply (gine2021mathematical, Theorem 3.5.1) on empirical processes and covering number. Therefore for all $\ell$:
$\mathbb{E}\left[\sqrt{n}\underset{f\in\mathcal{F}_{p}}{\sup}|\mathbb{P}_{n}^{\ell}(f)-\mathbb{E}_{X\sim\mu^{\ell}}[f(X)]|\right]\leq 8\sqrt{2}\mathbb{E}\left[\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2\mathcal{N}(\tau/2,\mathcal{F}_{p},\|\cdot\|_{\mathbb{L}^{2}(\mu_{n}^{\ell})})\right)}d\tau\right].$
From Lemma C.1 applied to the measure $\mu_{n}^{\ell}$, denoting $\tilde{K}=2(2R)^{p}$, we get
$\begin{split}\int_{0}^{\tilde{K}}\sqrt{\log(2\mathcal{N}(\tau/2,\mathcal{F}_{p},\|\cdot\|_{\mathbb{L}^{2}(\mu_{n}^{\ell})}))}d\tau&\leq\int_{0}^{\tilde{K}}\sqrt{\log\left(2\left[1+\frac{8pR^{p}}{\tau}\right]^{dN}\left[1+\frac{16pR^{p}}{\tau}\right]^{N}\right)}d\tau\\ &\leq\sqrt{N(d+1)}\int_{0}^{\tilde{K}}\sqrt{\log\left(2+\frac{32pR^{p}}{\tau}\right)}d\tau.\end{split}$